OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x - 9.

OpenStudy (solomonzelman):

subtract 4x from both sides subtract 2 from both sides

OpenStudy (anonymous):

i dont get it

OpenStudy (solomonzelman):

5x + 2 = 4x - 9 5x+2-4x=4x-9-4x x+2=-9 subtract 2 from both sides

OpenStudy (anonymous):

awwwwe ok